OpenStudy (anonymous):

Please help! I'm a little lost: The ratio of teachers to students at Washington Middle school is 1:15. If 900 student attend the middle school, how many teachers are there? Thx for helping ! : ]

OpenStudy (anonymous):

\[\frac{1}{15}=\frac{x}{900}\] \[x=\frac{900}{15}\]

OpenStudy (anonymous):

1 : 5 x : 900 1/5=x/900 900 = 5x x= 180

OpenStudy (anonymous):

i get \[\frac{900}{15}=60\]

OpenStudy (anonymous):

@ c0c0ple Uhh I think yuh wrote tha wrong thing

OpenStudy (anonymous):

Thanks!! ;]]